OpenStudy (barrelracing):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. a) -91 b) -33 c) -25 d) 35

OpenStudy (barrelracing):

@Michele_Laino

OpenStudy (anonymous):

First you're gonna do f(g(x)) so you plug in -5x+2 into the x's on the f(x) function \[f(g(x))=2(-5x+2)+1\] after that you just distribute and you get \[-10x+5\] now you just plug in x=3 into the equation and solve \[-10(3)+5\]\[-30+5\]\[=-25\]

OpenStudy (michele_laino):

by definition of composite function,, we can write this: \[f\left( {g\left( x \right)} \right) = 2g\left( x \right) + 1 = 2\left( { - 5x + 2} \right) + 1\]